OpenStudy (anonymous):

Simplify 9 - {x - (7 + x)}. 16 -2x + 2 -2x + 16 2x + 2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

meant b oops

OpenStudy (uri):

Show your work @dirtydan667

OpenStudy (anonymous):

i subtracted the like terms and got two but i dont know what to do now

OpenStudy (skullpatrol):

The first thing you need to know is the definition of subtraction a - b = a + (-b)

OpenStudy (skullpatrol):

9 - {x - (7 + x)} 9 - {x + (-7 + (-x))}

OpenStudy (anonymous):

i dont know dood

OpenStudy (anonymous):

can you explain it more

OpenStudy (skullpatrol):

To find the opposite of a sum or difference you change the sign of each term of the sum or difference -(a + b) = (-a) + (-b) or -(a - b) = (-a) + b

OpenStudy (skullpatrol):

- (7 + x) = -7 + (-x)

OpenStudy (anonymous):

im sorry my brain dosent work in the mornign

OpenStudy (skullpatrol):